OpenStudy (anonymous):

What is 32.5 of 1800

OpenStudy (♪chibiterasu):

\[\frac{ 1800 }{ 32.5 } = 55.38\] Rounded to the nearest hundreth.

OpenStudy (jdoe0001):

\(\large {\begin{array}{ccllll} quantity&percentage\\ \hline\\ 1800&100\\ 32.5&x \end{array}\implies \cfrac{1800}{32.5}=\cfrac{100}{x}}\)

OpenStudy (tkhunny):

32.5 = 3250% 32.5 * 1800 = 58500 We might need a better question.

OpenStudy (♪chibiterasu):

I think he meant 32.5%

OpenStudy (anonymous):

Thank you all